OpenStudy (anonymous):

what is the solution to the rational equation 3/2 - 1/x-5 = -2/2x-10

OpenStudy (anonymous):

\[{3(x-5)\over2(x-5)}-{1(2)\over(x-5)(2)}={-2\over(2)(x-5)(x-5)}\]

OpenStudy (anonymous):

i dont think thats right, cause i have four choices.. a.)x= -5 b.)x=-5/2 c.)no solution d.)x=5

OpenStudy (anonymous):

@optiquest

OpenStudy (anonymous):

do you not know what to do next?

OpenStudy (anonymous):

yeah, i'd have to solve it, i just get confused what do i do with the 2 in the denominator?

OpenStudy (anonymous):

the left side's denominators match, so subtract the two fractions. then work on isolating x to one side

OpenStudy (anonymous):

sorry ignore the first response. multiply all terms by (x)(x-5) and it will eliminate the fractions

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

would my answer be 5?

OpenStudy (anonymous):

@optiquest